7) If the corporate tax rate is 40% and the individual tax rate is 20% (and these are the only taxes), if a corporation has $50,000 of operating income, what is the after tax income for its owners? The tax rate is 15% between $0-$50,000 and 25% between $50,001-S 100,000. If you earn $60,000, then what is your average and marginal tax rates? A bank has quoted to you a monthly rate of 0.25% on a short-term loan. What is the stated annual rate of that loan? What is the future value of $8 in three years if the interest rate is 20% and the compounding is continuous? e What is the future value of a 3-year annuity due with $50 cash flows if the annually compounded interest rate is 6%? f)If we need to find the future value of a current cash flow and interest is compounded quarterly, how many periods of compounding occur in two and half years? e Whatis the diffrence between the present value of a 3-year ordinary annuity with a s20 cash flow and the present value of a 3-year annuity due with a $20 cash flow when the interest rate is 3.5%?

Please avoid using excel to solve the problem. Thanks!
Answer #1

(a) Corporate Tax Rate = 40 % and Individual Tax Rate = 25 %

Operating Income = $ 50000

After-Tax Net Income = Operating Income - Corporate Tax = 50000 x (1-0.4) = $ 30000

After-Tax Income Accruing to Firm owners will be again taxed at the individual tax rate of 25 %.

Therefore, final after-tax income of owners = 30000 x (1-0.25) = $ 22500

(b) Income = $ 60000, Tax Rate between $ 0 - 50000 = 15 % and Tax Rate between $ 50001 - 100000 = 25 %

Therefore, Total Tax Outlay = (50000 - 0) x 0.15 + (60000 - 50000) x 0.25 = $ 10000

Marginal Tax Rate = Limiting (Last) Tax Bracket Applicable = 25 %

Average Tax Rate = Tax Outlay / Income = 10000 / 60000 ~ 0.1667 or 16.67 %

(c) Quoted Monthly Rate = 0.25 %

Therefore, Stated (Simple Interest) Annual Rate = 0.25 x 12 = 3 %

(d) Applicable Interest Rate = 20%, Tenure = 3 years and Continuous Compounding, Initial Investment = $ 8

Therefore, Future Value = 8 x EXP[0.2 x 3] ~ $ 14.5769 ~ $ 14.58

NOTE: Please raise separate queries for solutions to the remaining sub-parts as one query is restricted to the solution of only one complete question (with a maximum of four sub-parts)